use the vectors v=-2i + j and w=4i -3j to find the following
4v – 3w .
Answer #1

given Vectors v=-2i+j W = 4i-3i calculate 4V Now 4 Eriti 40=-8itui calculate 3 il = 3(41-33) 121 -93 3 W 4 V-3W = [-si +4j) -